OpenStudy (anonymous):

Find the domain of the composite function f. g f(x)=2/x-3;g(x)=7/x

OpenStudy (anonymous):

f(x)=2/x-3;g(x)=7/x f(x)=2/(x-3) f(g)=2/(g-3) g(x)=7/x f(g(x))=2/(7/x-3)=2x/(7-3x) Domain all real numbers except 3,0, and 7/3?